OpenStudy (anonymous):

Use synthetic substitution to evaluate the polynomial for the given value x = 1/3. P(x) = 3x3 + 5x2 + 4x + 2 P(1/3) = ____ Enter only a number.

OpenStudy (anonymous):

Could you not just type that into a calculator?

OpenStudy (anonymous):

9.7

OpenStudy (anonymous):

^ knows how to use a calculator

OpenStudy (diyadiya):

Plug in 1/3 in place of x P(x) = 3x3 + 5x2 + 4x + 2 \[P(1/3)= 3(1/3)^3+5(1/3)^2+4(1/3)+2\]

OpenStudy (anonymous):

i keep gettin the same awnser and it say i got it wrong

OpenStudy (diyadiya):

\[=3 \left( \frac{1}{27} \right)+ 5\left( \frac{1}{9} \right)+ \frac{4}{3}+2\] \[= \frac{1}{9}+ \frac{5}{9}+ \frac{4}{3}+2\]

OpenStudy (diyadiya):

\[\frac{6}{9}+ \frac{4}{3}+2= \frac{2}{3}+ \frac{4}{3}+2 = \frac{6}{3}+2= ?\] I think now you can get the answer

OpenStudy (diyadiya):

got it?

OpenStudy (anonymous):

4?

OpenStudy (amistre64):

what is synth sub?